OpenStudy (anonymous):

Factor the algebraic expression below in terms of a single trigonometric function. csc^2x - 1 its supposd to be in two bionmials. But how?

OpenStudy (anonymous):

probably not because the answer isnt cot^2x

OpenStudy (anonymous):

so you got cot^2 x answer !

OpenStudy (anonymous):

yeah but its not right

OpenStudy (anonymous):

how do u know ?

OpenStudy (anonymous):

my teacher said so

OpenStudy (anonymous):

cot^2x because- if you put csc^2x-1 in terms of sin and cos, you get... 1/(sin^2x)−1 continue putting the numerator in terms of sin and cos... =(sin^2x+cos^2x)/(sin^2x)−1 =1+(cos^2x)/(sin^2x)−1 The ones cancel out leaving... cos^2x/sin^2x which simplifies to... cot^2x ESSAY FEEDBACK Hi Amanda, You need to factor these into 2 binomials.

OpenStudy (anonymous):

okay here's the hint the question says two binomials !

OpenStudy (anonymous):

I know :/ I dont get it

OpenStudy (anonymous):

and again it says single trigonometric function !

OpenStudy (anonymous):

... cosx/sinx?

OpenStudy (anonymous):

add a two in front ov each x

OpenStudy (anonymous):

nope !

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

you'll get cos^2x/ sin^2x

OpenStudy (anonymous):

thats what i said...thats not right?

OpenStudy (anonymous):

which is binomial but still single and are two !

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

ur question says one trigonometric function but having two binomial terms

OpenStudy (anonymous):

I dont see how to do it then :/

OpenStudy (anonymous):

there's nothing to do !

OpenStudy (anonymous):

you already got the answer !

OpenStudy (anonymous):

b..b...but?

OpenStudy (anonymous):

it said it was wrong

OpenStudy (anonymous):

\[\cos^2x/ \sin^2x \]

OpenStudy (anonymous):

oh...so I just had to break it down more?

OpenStudy (anonymous):

yah!

OpenStudy (anonymous):

so you got it !

OpenStudy (anonymous):

Thanks a heap

OpenStudy (anonymous):

ur welcome